# The Unbearable Slowness of Being

### v1

@@ B.2 Language
Of the entire human motor system, our vocal tracts are the most specialized in communicating information. Analyses of speech recordings in 17 languages peg human speech at a universal transmission rate at $\sim$ 39 bits/s [coupe_different_2019]. Here, the information rate is the product of information density and speech rate, but the two are inversely correlated because of the speed-accuracy trade-off: fast speech tends to consist of more frequently-used but less informative words [mackay_problems_1982, cohen_priva_not_2017]. Speakers tend to slow down when they introduce contextually unexpected vocabulary. In addition, this number does not take into account much of the redundancy in language: as acknowledged in the methods section of the article, the true information rate is lower. Still, the estimate is in the same order of magnitude as our ballpark number.
@@
Even in cases of minimal movement, the speed of language processing remains in the same range. During reading, for example, one fixation typically spans 200-250 ms and covers 7-9 letters [rayner_eye_1998]. This yields an average reading speed of 28 - 45 English characters per second or bits/s. Similar to fast speech, a speed reader can skim through the text and triple their rate up to 700 words per minute or $\sim$ 60 characters/s, but comprehension is compromised [masson_conceptual_1983]. If one includes this error rate, the rate of reading remains considerably below 60 bits/s.
B.3 Perception
@@
What about innate or quickly learned responses to visual stimuli? The extensive research on human reaction time explored the upper limit of such responses. In a choice-reaction experiment by hick_rate_1952, participants made movements corresponding to the flashes of two to ten light bulbs. The light bulbs flickered in random order; therefore, to press the correct corresponding key was to make one out of 2 - 10 equally probable choices. As the number of lamps increased, the reaction time – at hundreds of milliseconds – was found to increase logarithmically. hyman_stimulus_1953 repeated the experiment in units of bits, which resulted in a linear relation and a constant slope – the information of the choices divided by the reaction time, or the information rate of the reaction. Over thousands of trials, “the rate of gain of information is, on the average, constant with respect to time”, at 5 bits/s. Different experimental designs reached a similar conclusion [klemmer_rate_1969], with information rates ranging from a few to ten bits/s.
@@
In cases where the need for motor output is eliminated, the rate might be slightly higher, at 30-50 bits/s for object recognition [sziklai_studies_1956]. In this study, each participant viewed multiple line drawings of objects that were shown simultaneously. When a frame of 16 objects was shown for 500 ms, all 50 participants recognized at least one independent object; the majority identified two, and only one individual managed to name three in a trial. sziklai_studies_1956 then modeled object recognition as selecting from a repertoire of a thousand basic picturable nouns – based on the vocabulary an adult English speaker should know. Thus each named item corresponds to 20 bits of information. Divide 20 bits by 500 ms and we get 40 bit/s on average.
@@
Behavior/Activity Time scale Information rate (bits/s) References
Binary digit memorization 5 min 4.9  [internationalassociationofmemory_5_2024]
Blindfolded speedcubing 12.78 sec 11.8  [guinness_world_records_limited_fastest_2023]
Choice-reaction experiment min $\sim$ 5 [hick_rate_1952, hyman_stimulus_1953, klemmer_rate_1969]
Listening comprehension (English) min - hr $\sim$13 [williams_guidelines_1998]
Object recognition 1/2 sec 30 - 50 [sziklai_studies_1956]
Optimal performance in laboratory motor tasks $\sim$15 sec 10 - 12 [fitts_information_1954, fitts_information_1964]
Reading (English) min 28 - 45 [rayner_eye_1998]
Speech in 17 languages < 1 min 39 [coupe_different_2019]
Speed card 12.74 sec 17.7 [internationalassociationofmemory_speed_2024]
StarCraft (e-athlete) min 10  [guinness_world_records_limited_most_2023]
Tetris (Rank S) min $<6$  [tenchi_tetrio_2021]
Typing (English) min - hr $10$  [dhakal_observations_2018, shannon_prediction_1951]
Table 1: The information rate of human behaviors
Appendix C The information capacity of neurons
@@ C.1 Photoreceptors
C.2 Spiking neurons
The typical spiking neuron can be viewed as a point-process channel: It converts a continuous time-varying input, like the synaptic current $s(t)$, into a discrete train of spikes at the output, $r=\{t_{i}\}$. Several approaches have been developed to estimate the resulting information rate [borst_information_1999]. A popular one is by “decoding the spike train”: Propose a model to estimate the stimulus $s_{\rm{est}}(t)$ from the spike train; the simplest such model is linear filtering of the spike train with some kernel $k(t)$,
@@ C.2 Spiking neurons
From an experiment that measures both $s$ and $r$, optimize the decoding kernel $k(t)$ so as to minimize the difference between $s(t)$ and $s_{\rm{est}}(t)$. Finally measure the mutual information between $s(t)$ and $s_{\rm{est}}(t)$. These are both continuous functions, so the Shannon-Hartley theorem applies, though one needs to be careful along the way in managing lower and upper bounds so as to maintain a conservative estimate. For many neuron types, the information capacity of the spike train has been measured at 1-4 bits/spike [borst_information_1999].
C.3 Capacity vs information rate
@@
In normal function, both people and single neurons are rarely driven at the maximal rate specified by the capacity. People reach their capacity under special professional conditions (typist) or the pressure of competition (Starcraft player). Sensory neurons are found to reach their capacity preferentially during biologically important tasks, like the discrimination of mating calls (frogs), or the real-time control of aerobatics (flies) [bialek_reading_1991, rieke_naturalistic_1995].
C.4 Some caveats
We adopt the use of information theoretic measures here, because they allow a comparison of behavioral performance in neurons, networks, entire organisms, and machines using the same units of bits per second. In simple terms, a behaving agent chooses one among many possible actions given a set of inputs, and the information rate measures how much that range of actions gets constrained by the decisions. Information theory is not the only statistical framework one could employ, and we don’t consider it a magic bullet [shannon_bandwagon_1956].
@@ Acknowledgments
For discussions and comments on the manuscript we thank Frederick Eberhardt, Michelle Effros, Florian Engert, Mehrdad Jazayeri, Christof Koch, Ueli Rutishauser, and Anthony Zador.

### v2

@@ A.2 Language
Of the entire human motor system, our vocal tracts are the most specialized in communicating information. Analyses of speech recordings in 17 languages peg human speech at a universal transmission rate of 39 bits/s coupe_different_2019. Here, the information rate is the product of information density and speech rate, but the two are inversely correlated because of the speed-accuracy trade-off: fast speech tends to consist of more frequently-used but less informative words mackay_problems_1982, cohen_priva_not_2017. Speakers tend to slow down when they introduce contextually unexpected vocabulary. In addition, this number does not take into account much of the redundancy in language: as acknowledged in the methods section of the article, the true information rate is lower. Still, the estimate is in the same order of magnitude as our ballpark number.
@@
Even in cases of minimal movement, the speed of language processing remains in the same range. During reading, for example, one fixation typically lasts 200-250 ms and covers 7-9 letters rayner_eye_1998. This yields an average reading speed of 28 - 45 English characters per second or bits/s. Similar to fast speech, a speed reader can skim through the text and triple their rate up to 700 words per minute or $\sim$ 60 characters/s, but comprehension is compromised masson_conceptual_1983. If one includes this error rate, the rate of reading remains considerably below 60 bits/s.
A.3 Perception
@@
What about innate or quickly learned responses to visual stimuli? The extensive research on human reaction time explored the upper limit of such responses. In a choice-reaction experiment by Hick hick_rate_1952, participants made movements corresponding to the flashes of two to ten light bulbs. The light bulbs flickered in random order; therefore, to press the correct corresponding key was to make one out of 2 - 10 equally probable choices. As the number of lamps increased, the reaction time – at hundreds of milliseconds – was found to increase logarithmically. Hyman hyman_stimulus_1953 repeated the experiment in units of bits, which resulted in a linear relation and a constant slope – the information of the choices divided by the reaction time, or the information rate of the reaction. Over thousands of trials, “the rate of gain of information is, on the average, constant with respect to time”, at 5 bits/s. Different experimental designs reached a similar conclusion klemmer_rate_1969, with information rates ranging from a few to ten bits/s.
@@
In cases where the need for motor output is eliminated, the rate might be slightly higher, at 30-50 bits/s for object recognition sziklai_studies_1956. In this study, each participant viewed multiple line drawings of objects that were shown simultaneously. When a frame of 16 objects was shown for 500 ms, all 50 participants recognized at least one independent object; the majority identified two, and only one individual managed to name three in a trial. Sziklai sziklai_studies_1956 then modeled object recognition as selecting from a repertoire of a thousand basic picturable nouns – based on the vocabulary an adult English speaker should know. Thus each named item corresponds to 20 bits of information. Divide 20 bits by 500 ms and we get 40 bit/s on average.
@@
Appendix B The information capacity of neurons
@@ B.1 Photoreceptors
B.2 Spiking neurons
The typical spiking neuron can be viewed as a point-process channel: It converts a continuous time-varying input, like the synaptic current $s(t)$, into a discrete train of spikes at the output, $r=\{t_{i}\}$. Several approaches have been developed to estimate the resulting information rate borst_information_1999. A popular one is by “decoding the spike train”: Propose a model to estimate the stimulus $s_{\rm{est}}(t)$ from the spike train; the simplest such model is linear filtering of the spike train with some kernel $k(t)$,
@@ B.2 Spiking neurons
From an experiment that measures both $s$ and $r$, optimize the decoding kernel $k(t)$ so as to minimize the difference between $s(t)$ and $s_{\rm{est}}(t)$. Finally measure the mutual information between $s(t)$ and $s_{\rm{est}}(t)$. These are both continuous functions, so the Shannon-Hartley theorem applies, though one needs to be careful along the way in managing lower and upper bounds so as to maintain a conservative estimate. For many neuron types, the information capacity of the spike train has been measured at 1-4 bits/spike borst_information_1999.
B.3 Capacity vs information rate
@@
In normal function, both people and single neurons are rarely driven at the maximal rate specified by the capacity. People reach their capacity under special professional conditions (typist) or the pressure of competition (Starcraft player). Sensory neurons are found to reach their capacity preferentially during biologically important tasks, like the discrimination of mating calls (frogs), or the real-time control of aerobatics (flies) bialek_reading_1991, rieke_naturalistic_1995.
@@ Acknowledgments
For discussions and comments on the early stage of the manuscript, we thank Frederick Eberhardt, Michelle Effros, Florian Engert, Mehrdad Jazayeri, Christof Koch, Ueli Rutishauser, and Anthony Zador. We also thank alphaXiv for providing a platform for open discussions on our preprint, where the comments received were intriguing and helpful for the revision.
@@
Declaration of Interests
The authors declare no competing interests.